Birr 6‚500 were divided equally among a certain number of persons. Had there been 15 more persons, each would have got Birr 30 less. Find the originals number of persons.


1
Expert's answer
2021-04-09T11:59:06-0400

"\frac{6500}{x}=\frac{6500}{x+15}+30,"

"x^2+15x-3250=0,"

"x=\frac{-15±\sqrt{225+13000}}{2}=\frac{-15±115}{2},"

"x>0,"

"x=50~(persons)."